sin inverse root{2-root3)/4}

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
t = sin^{-1}\sqrt{\frac{2-\sqrt{3}}{4}}
sint = sinsin^{-1}\sqrt{\frac{2-\sqrt{3}}{4}}
sint = \sqrt{\frac{2-\sqrt{3}}{4}}
sin^{2}t = {\frac{2-\sqrt{3}}{4}}
\frac{1-cos2t}{2} = {\frac{2-\sqrt{3}}{4}}
2-2cos2t = {\frac{2-\sqrt{3}}{1}}
cos2t=\frac{\sqrt{3}}{2}
2t = \frac{\pi }{6}
t = \frac{\pi }{12}
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty